Solve the system by substitution: x + y = 6 and x - y = 10. Show work.
A
(-8, -1)
B
(1, -3)
C
(7, -1)
D
(8, -2)

Answer by lynnlo(4176)

ANSWER:
D
PROBLEM:
X+Y=6
X-Y=10
STEP(1)=====ADD THE TWO EQUATIONS TOGETHER TO ELIMINATE "Y"
2X=16
STEP(2)=====FIND "X"
X=8
STEP(3)=SUBSTITUTE THE VALUE FOR "X" INTO THE EQUATION
TO SOLVE FOR"Y"
(8)+Y=6
Y=-2